ROTATIONAL ANALYSIS OF THE VIBRATIONAL GROUND STATE OF ACETONE, CH_{3}COCH_{3}

\underline{P. GRONER}, Department of Chemistry, University of Missouri - Kansas City, Kansas City, MO 64110; S. ALBERT, E. HERBST, F. C. DE LUCIA, Department of Physics, Ohio State University, Columbus, OH 43210; and F. J. LOVAS, National Institute of Standards and Technology, Optical Technology Division, Gaithersburg, MD 20899.

An effective rotational Hamiltonian^{a} was used to analyze rotational transitions in the vibrational ground state of acetone. Microwave and mm-wave measurements from the literature^{b} were combined with new measurements between 260 and 350 GHz and with new FT microwave frequencies in a global fit of all four torsional substates. Over 500 frequencies between 8 and 350 GHz were fit for transitions involving energy levels with J up to 34 and K_{a} up to 10. In one (preliminary) fit, 31 spectroscopic parameters were used to fit 555 frequencies to a dimensionless standard deviation of 1.25. The parameters determined in the least-squares fit were: ρ = 0.0621754(64), β = 25.8160(73) deg., parameters equivalent to the rotational, quartic and sextic distortion constants, the internal energy tunneling parameters ϵ_{01} = −763.436(83) MHz, ϵ_{1−1} = 0.1057(62) MHz, ϵ_{11} = 0.991(34) MHz, ϵ_{02} = 0.585(14) MHz and ten tunneling constants related to the rotational and distortion constants. This particular fit is not quite satisfactory because a number of transitions, including low K_{c} R-transitions, had to be excluded from the fit.

aP. Groner, J. Chem. Phys. 107, 4483 (1997).

bR. Peter and H. Dreizler, Z. Naturforsch. A 20, 301 (1965); J. M. Vacherand, B. P. Van Eijck, J. Burie and J. Demaison, J. Mol. Spectrosc. (1986) 118, 355; F. Oldag and D. H. Sutter, Z. Naturforsch. A (1992) 478, 527 (1992).